# Stochastic gravitational-wave background search using data from five pulsar timing arrays

### v1

@@ Abstract
Using public data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB). We use pulse-time-of-arrival data from 121 pulsars, which is more informative than previous searches, carried out separately by the individual PTA collaborations using only their own data. For pulsars observed by more than one PTA, we employ a new “direct combination” method to merge their astrophysical models and data. This avoids the challenge of reconciling the different PTA timing models to obtain a single “best” model. A central result of our analysis is posterior probability distributions for the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency. While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and through noise-marginalized false-alarm probabilities ($p$-values) for three different detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection. We also reconstruct the inter-pulsar timing-residual correlation, as a function of the angle $\theta$ between the pulsar lines of sight. This is consistent with the curve predicted by Hellings and Downs (HD).
@@
Introduction – Pulsar timing arrays (PTAs) aim to detect low frequency gravitational waves (GWs), by searching for the effects that such waves induce on the arrival times of pulsar pulses. These GWs create “timing residuals” which are correlated among different pulsars [1, 2]. Decades-long datasets with submicrosecond timing precision yield GW amplitude (strain) sensitivities of order $A_{gw}\sim 10^{-15}$ at nHz frequencies.
@@
Orbiting pairs of supermassive black holes (SMBHs) are one possible source of GWs. If the Universe contains many such pairs, then the incoherent superposition of their GW emissions would create a stochastic GW background (SGWB) [3, 4, 5] with a power-law spectrum in the nHz band.
@@
Different PTA collaborations have reported varying levels of evidence for a nHz SGWB. The Parkes PTA (PPTA [6]) observed 30 pulsars over 18 years, concluding that the data showed “no support for or against”. The European and Indian PTAs (EPTA and InPTA [7]) observed 25 pulsars, and reported “marginal evidence” in the 25-year dataset and “evidence” when only the last 10 years of data were analyzed. The Chinese PTA (CPTA [8]) observed 57 pulsars over three years, and reported “some evidence”. The North American Nanohertz Observatory for Gravitational Waves (NG [9]) observed 67 pulsars over 15 years and reported “compelling evidence”. Most recently, the MeerKAT PTA (MPTA [10]) observed 83 pulsars over 4.5 years, and, depending on assumptions about pulsar noise processes, reported either “high significance” or “no significance” (also see [11]). None of these pass the threshold of $5\sigma$ significance required to claim detection [12].
@@
Figure 1: A Mollweide projection showing sky locations of the 121 pulsars included in our analysis. Colors indicate how many PTAs contributed data for each pulsar.
@@
All of these PTAs (apart from CPTA) have released their data to the public. Using this full public dataset increases the total number of distinct pulsars to $N=121$, as shown in Fig. 1. PTAs search for an SGWB by correlating pairs of pulsars, whose number scales as $N^{2}$. Furthermore, since most pulsars are observed by more than one collaboration, there are also more data samples per pulsar. Hence, analysis of the full dataset should significantly increase the GW strain sensitivity and the detection confidence [13, 14, 15, 2].
@@
We analyze this five-PTA dataset [16, 17, 18, 19, 20] using a recently proposed “direct combination” method [21, 22]. This allows us to combine the data with only minor modifications, as described below. Tests of this method, carried out with the International Pulsar Timing Array (IPTA) Data Release 2 (DR2) combined data set from 2019 [14], suggest that it produces results that are identical to traditional (but more time-consuming) data combination techniques.
@@
Aside from this, our analysis is standard. We use the same tools and methods as the individual PTAs, as implemented in Enterprise [23] and in the next-generation PTA data-analysis package Discovery [24]. Both are based on parameterized likelihood models and employ Markov chain Monte Carlo methods. These produce posterior samples (here called “posteriors”) that approximate the inferred posterior probability distributions of the model parameters.
@@
Methods, Models, and Conventions – Because we use standard methods, this paper is brief. When we follow the PTA analyses described above [6, 7, 8, 9, 10], we give only a brief description and cite the relevant literature. If we do something different, then more details are provided.
@@
– The null hypothesis is called “common uncorrelated red noise” and is denoted by CURN. Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$. Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis. The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the HD correlations [27]. It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms. Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR. Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2]. CURN is considered a more conservative null hypothesis than that of “no GWs”.
Using Discovery [24] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models. These are sets of parameter values that are most the consistent with the data and priors; they characterize the joint posterior probability distributions of the model parameters.
For this work, the most important model parameters are the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the SGWB power-law power spectrum. To define these, let ${\rm{d}}\rho_{gw}$ denote the SGWB energy density in the frequency range $f,f+{\rm{d}}f]$ and let $\rho_{cr}=3c^{2}H_{0}^{2}/8\pi G\simeq 7.7\times 10^{-10}\ \mathrm{J}/\mathrm{m}^{3}$ denote the critical (closure) energy density of the Universe. Here, $c$ is the speed of light, $G$ is Newton’s gravitational constant, and $H_{0}\simeq 67.4~\mathrm{km}/\mathrm{s\,Mpc}\simeq 2.2\times 10^{-18}\,\mathrm{Hz}$ is the current Hubble expansion rate. Then
@@
where $f_{r}=1/\mathrm{yr}\simeq 31.7\,\mathrm{nHz}$ is a standard reference frequency. Together with our previous assumptions, this spectral model (assumed to apply in the range $2<f/\mathrm{nHz}<200$) fully characterizes the SGWB.
One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$. We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$. The latter would be expected from a population of SMBHs in circular binaries, provided that GWs dominate the energy loss at nHz frequencies.
@@
Each of the posteriors corresponds to a different model universe, with its own amplitudes and exponents for the different noise sources, and thus with its own covariance matrices $\bf C$ and $\bf C_{0}$. This means that the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next. Thus, to assess detection confidence, we present the posterior distributions of these $p$-values. In the literature, its mean is called the “noise-marginalized false-alarm probability” [28] or “posterior predictive $p$-value” [29].
Some past work uses ad-hoc methods for assessing the false-alarm probability $p$. “Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”. The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [30, 31]. The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [31]. However, neither of these methods properly reproduces the null distribution of $D$ [32]. For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
We determine the false-alarm probability $p$ using a correct method [33, 34, 32, 27]. Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [33, 32, 27]; the probability that it exceeds the observed value in one of our posterior samples is found analytically.
@@
Analysis of the five-PTA dataset – The data from the five PTAs come from 121 millisecond pulsars, as shown in Fig. 1. These consist of the second EPTA data release (10-year subset) [35, 16], the first InPTA data release [36, 17], the MPTA 4.5-yr data release [37, 18], the NG 15yr data release [38, 19], and the third PPTA data release [39, 20]. Each PTA’s data consists of one .par file and one or more .tim files, per pulsar.
We merge these using the direct combination method [21, 22]. The .tim files, containing pulse times-of-arrival (TOAs) and their uncertainties, are used “as is”. The .par files, containing the astrophysical models, and some detector-specific parameters, are modified (in effect) as follows, one pulsar at a time. (This is also the first step of “manual” data combination.)
@@
    For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$. (InPTA is not listed because its pulsars are observed by other PTAs.) Any other PTAs that have data for this pulsar are called “target PTAs”.
@@
    Any .par file that uses the TCB time standard is converted to TDB using Tempo2 [40] tools.
@@
    If the reference .par file has a DMX model for DM fluctuations (NG and InPTA only), then it is replaced by a Gaussian process (DMGP) model. The parameter DM0 is initialized to the mean DMX value, and DM1 and DM2 are initialized to zero.
@@
    If the reference .par file has a DMGP model, then the DM0 value is retained, and DM1 and DM2 are initialized to zero.
  6. 6.
    The astrophysical model $\beta$ [1, Eq. 7.2], including sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values, is copied from the reference .par file to the target .par files.
Detector-specific parameters in the .par files, for example, “JUMP” for time offsets and “FDx” for narrow-band frequency offsets, are unchanged. This allows for template model shape differences between PTAs.
For four pulsars, ad-hoc “tricks” are used to maintain phase connection between the .tim file TOAs. For the binary pulsar PSR J2241-5236 (MPTA, PPTA), we retain the MPTA 11-parameter model and the PPTA 18-parameter model for the orbital frequency evolution. The other parameters follow 6 above. For PSR J0610-2100 (MPTA, NG) and PSR J1024-0719 (EPTA, MPTA, NG, PPTA), we keep the DMX model for the NG data and the DMGP model for the remaining PTAs. For PSR J1804-2858 (MPTA), we retain the values of DM1 and DM2 rather than initializing them to zero.
The direct combination method [21, 22] works because the only significant change it makes is to the parameters $\beta$, item 6 above. While their exact values are unknown, every pulsar has unique values for these parameters, for example, the sky position. Those enter the analysis via a nominal value $\beta_{0}$ [1, Eq. 7.2] and via a matrix $\mathbf{M}$} [1, Eq. 7.4], which is the rate of change of the pulse arrival-time with respect to variations $\epsilon$ in $\beta$. In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} with the same $\beta_{0}$ for a given pulsar in all PTAs introduces negligible errors. The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq. 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs. Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same for all PTAs.
The next analysis step is single-pulsar noise modeling. The goal is to determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq. 7.27]. This characterizes the timing residuals fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eq. 7.2,7.26]. Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
Single-pulsar noise modeling determines two or three parameters per pulsar and backend. The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents. NG, PPTA, MPTA, and seven EPTA backends use the parameters $EFAC$, EQUADEQUAD, and $ECORR$ [38, Eq. 2]. InPTA and the remaining EPTA backends only use $EFAC$ and $EQUAD$ [41].
@@
Table 1: Priors for the WN parameters, and for the amplitudes/exponents of the power-law RN, DM and SGWB noise spectra. “S” denotes single pulsar, and “F” denotes full PTA.
The WN parameters for a given pulsar are obtained from all of its TOA data (meaning, all .tim files). These are analyzed using JAXopt [42], with priors taken from the Table 1 rows with an “S” in the final column; the amplitude and exponent priors are used for all power laws. For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model; for $\gamma>1$, the $f\to\infty$ divergences are absent [43, 44]. Note that $EQUAD$ has the units used in Tempo [45], Tempo2 [40] and PINT [46], and is often written as $T2EQUAD$. The output is a “WN dictionary” of best-fit $EFAC$, EQUADEQUAD and (potentially) $ECORR$ values, which are used in subsequent analysis to obtain $\mathbf{N}$} [38, Eq. 2].
For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA-analysis values. This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters.
We drop TOAs for any backends which have $EFAC<0.1$ or $EFAC>10$, since this indicates that the backend or template matching was unstable, or that RFI or calibration or instrumental issues were present. This removes $0.17\%$ of the TOAs. Our final data for analysis comes from 25 EPTA pulsars, 13 (of 14) InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 31 (of 32) PPTA pulsars. (Data from InPTA pulsar J1022+1001 are dropped because all the $EFAC$s are larger than 10. Data from PPTA pulsar PSR J1741+1351 are dropped because there were only 16 observations.)
@@
Within the Discovery framework, we use the priors listed in Tab. 1 with an “F” in the final column. As described above, we consider two priors for the SGWB power-law power-spectrum exponent $\gamma_{gw}$. One is the uniform range given in Tab. 1, the other is fixed to $\gamma_{gw}=13/3$. Our analysis uses the Hamiltonian No-U-Turn Sampler (NUTS) [48, 49], as implemented in NumPyro [50, 51, 52]. NUTS is the default for Discovery [24], which provides the log-likelihood gradients needed by Hamiltonian samplers. It is much faster than the PTMCMC sampler [53] used by the individual PTAs when they analyzed their own data.
We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number of posteriors under the CURN hypothesis. Each gives the amplitudes and exponents of the power-law models describing the pulsar (red and DM) noise, and the SGWB. We use these to evaluate the frequentist detection statistics and their corresponding false-alarm probabilities ($p$-values) for each posterior sample.
@@
Results of the five-PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1). These are shown in Fig. 2. The amplitude posterior is sharply peaked away from zero, which is consistent with (or evidence for) the presence of an SGWB. As expected, the posteriors are very similar for the two hypotheses. This is because, if there are $\lesssim 1500$ pulsars (see [54, Fig. 10]), most of the SGWB evidence arises from autocorrelations rather than cross-correlations—–and by construction, CURN includes these autocorrelations.
@@
CURN with free $\gamma_{gw}\qquad$ $14.24^{+0.11}_{-0.11}\qquad$ $3.30^{+0.28}_{-0.27}$

### v2

@@ Abstract
Using public pulse time-of-arrival data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB). This analysis is more sensitive than previous individual PTA searches because the combined 121-pulsar dataset is about four times larger than any single PTA’s. For pulsars observed by multiple PTAs, we employ a new “direct combination” method to merge their astrophysical models and data. This avoids the challenge of reconciling different PTA timing models to obtain a single “best” model. A central result of our analysis is the posterior distribution of the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency. While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection. The inter-pulsar timing-residual correlation, reconstructed as a function of angle $\theta$ between the pulsar lines of sight, matches the Hellings and Downs (HD) prediction.
@@
Introduction – Pulsar timing arrays (PTAs) aim to detect low-frequency gravitational waves (GWs) by measuring the effects these waves induce on the arrival times of pulsar pulses. These GWs create “timing residuals” which are correlated between pulsars [1, 2]. Decades-long datasets with submicrosecond timing precision yield GW amplitude (strain) sensitivities of order $A_{gw}\sim 10^{-15}$ at nHz frequencies.
@@
Supermassive black hole binaries (SMBHBs) are one source of GWs. If the Universe contains many of them, the incoherent superposition of their GW emissions would create a stochastic GW background (SGWB) [3, 4, 5] with a power-law spectrum in the nHz band.
@@
Different PTA collaborations have reported varying levels of evidence for a nHz SGWB. The Parkes PTA (PPTA [6]) observed 30 pulsars over 18 yr, concluding that the data show “no support for or against”. The European and Indian PTAs (EPTA and InPTA [7]) observed 25 pulsars and reported “marginal evidence” in the 25-yr dataset and “evidence” when only the last 10 yr were analyzed. The Chinese PTA (CPTA [8]) observed 57 pulsars over 3 yr, and reported “some evidence”. The North American Nanohertz Observatory for Gravitational Waves (NG [9]) observed 67 pulsars over 15 yr and reported “compelling evidence”. Most recently, the MeerKAT PTA (MPTA [10]) observed 83 pulsars over 4.5 yr, and, depending on assumptions about pulsar noise processes, reported either “high significance” or “no significance” (see also [11]). None reach the conventional $5\sigma$ significance threshold required to claim a detection [12].
@@
Figure 1: A Mollweide projection showing sky locations of the 121 pulsars included in our analysis. Colors indicate how many PTAs provided data for that pulsar.
@@
All of these PTAs (apart from CPTA) have released their data. Using this full public dataset increases the total number of distinct pulsars to $N_{p}=121$ as shown in Fig. 1. PTAs search for an SGWB by correlating pairs of pulsars, whose number scales as $N_{p}^{2}$. Furthermore, since many pulsars are observed by more than one collaboration (on average, 1.8) there are also more data samples per pulsar. Hence, analysis of the combined dataset should significantly increase the GW strain sensitivity and the detection confidence [13, 14, 15, 2].
@@
We analyze this five PTA dataset [16, 17, 18, 19, 20] using a recently proposed “direct combination” method [21, 22]. This allows us to combine the data with minor modifications, as described below. Tests of this method [21], carried out with the International Pulsar Timing Array [23] (IPTA) Data Release 2 (DR2) combined dataset from 2019 [14], suggest that it produces results identical to traditional (but more time-consuming) data combination techniques.
@@
Aside from this, our analysis is standard. We use the same tools and methods as the individual PTAs, as implemented in Enterprise [24] and in the next-generation PTA data-analysis package Discovery [25]. Both are based on parameterized likelihood models, and employ Markov chain Monte Carlo (MCMC) methods to produce posterior samples (here called “posteriors”) which approximate the inferred posterior probability distributions of the model parameters. Detection confidence is assessed using both Bayesian and frequentist methods.
Note: additional tables and plots may be found in the Supplementary Material (SM).
@@
Methods, Models, and Conventions – Because we use standard methods, this paper is brief. When we follow the PTA analyses described above [6, 7, 8, 9, 10], we give only a cursory description and cite the relevant literature. If we do something different, then more details are provided.
@@
– The null hypothesis is called “common uncorrelated red noise” and is denoted by CURN. Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$. Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis. The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the name “uncorrelated”. It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms. Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR. Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2]. CURN is a more “conservative” null hypothesis than that of “no GWs”.
Using Discovery [25] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models. These are sets of inferred parameter values most consistent with the data and priors, thereby characterizing the posterior distributions of the model parameters.
The most important model parameters are $A_{gw}$ and $\gamma_{gw}$, which control the power-law amplitude and exponent for the SGWB power spectrum. To define these, let ${\rm{d}}\rho_{gw}$ denote the SGWB energy density in the frequency range $f,f+{\rm{d}}f]$ and $\rho_{cr}=3c^{2}H_{0}^{2}/8\pi G\simeq 7.7\times 10^{-10}\ \mathrm{J}/\mathrm{m}^{3}$ denote the critical (closure) energy density of the Universe. Here, $c$ is the speed of light, $G$ is Newton’s gravitational constant, and $H_{0}\simeq 67.4~\mathrm{km}/\mathrm{s\,Mpc}\simeq 2.2\times 10^{-18}\,\mathrm{Hz}$ is the current Hubble expansion rate. Then
@@
where $f_{r}=1/\mathrm{yr}\simeq 31.7\,\mathrm{nHz}$ is a standard reference frequency. Together with our previous assumptions, this spectral model fully characterizes the SGWB over the 14 frequency bins of our analysis $2\lesssim f/\mathrm{nHz}\lesssim 30$.
One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$. We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$. The latter would be expected from a population of SMBHBs in circular orbits, provided that GWs dominate the energy loss at nHz frequencies.
@@
Each posterior corresponds to a different model universe, with its own amplitudes/exponents for the different noise sources, and thus its own covariance matrices $\bf C$ and $\bf C_{0}$. Hence, the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next. To assess detection confidence, we present the posterior distributions of these $p$-values. In the literature, its mean is called the “posterior predictive $p$-value” [29], similar in spirit to the “noise-marginalized optimal statistic” [30].
Some past work uses ad hoc methods for assessing the false-alarm probability $p$. “Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”. The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [31, 32]. The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [32]. However, neither of these methods properly reproduces the null distribution of $D$ [33]. For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
We determine the false-alarm probability $p$ using a statistically rigorous method [34, 35, 33, 28]. Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [34, 33, 28]; the probability that it exceeds the observed value in each of our posterior samples is found analytically.
@@
Analysis of the five PTA dataset – The data from the five PTAs come from 121 millisecond pulsars, as shown in Fig. 1. These consist of the second EPTA data release (10-yr subset) [36, 16], the first InPTA data release [37, 17], the MPTA 4.5-yr data release [38, 18], the NG 15-yr data release [39, 19], and the third PPTA data release [40, 20]. Each PTA’s public open dataset has one .par file and one or more .tim files per pulsar.
The data we analyze consist of $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ times of arrival (TOAs): the complete set of observations from 25 EPTA pulsars, 14 InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 32 PPTA pulsars. We merge these using the direct combination method [21, 22]. If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties. The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
@@
    For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$. InPTA is not listed because its pulsars are observed by other PTAs. Other PTAs that have data for this pulsar are called “target PTAs”.
@@
    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [41] tools. See SM Sec. III.
@@
    The reference .par file DM model is replaced with a constant-in-time DM model: DM0 is initialized to the .par file DM value at time DMEPOCH, and DM1 and DM2 are initialized to zero.
@@
    A nominal astrophysical model $\beta_{0}$ [1, Eq. 7.2] (sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values) is copied from the reference .par file to the target .par files.
Detector-specific parameters in the .par files, for example “FDx” for narrow-band frequency offsets, are unchanged. One additional adjustable phase shift parameter (“JUMP”) per target PTA is introduced to account for pulse profile template shape differences.
The direct combination method [21, 22] adopts one set of parameter values $\beta_{0}$ for each pulsar, item 5 above. It works because each pulsar has exact but unknown values $\beta$ for these (for example, for the sky position). They enter the analysis via the nominal value $\beta_{0}$ [1, Eq. 7.2] and via a “design matrix” $\mathbf{M}$} [1, Eq. 7.4], which is the rate of change of the pulse arrival-time with respect to variations in $\beta$, evaluated at $\beta_{0}$. In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} at the same nominal $\beta_{0}$ for a given pulsar introduces negligible errors. The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq. 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs. Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same across PTAs. We therefore adopt the physically motivated DMGP model [42], used by all PTAs except NG and InPTA, which use the DMX model. The design matrix $\mathbf{M}$} allows the effects of DM fluctuations on the timing residuals to be computed from the fixed-in-time DM timing residuals.
The next analysis step is single-pulsar noise modeling. These determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq. 7.27]. This characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs. 7.2 and 7.26], modeled with 30 and 100 frequency bins respectively. Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
Single-pulsar noise modeling determines two or three parameters per pulsar and backend. The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents. NG, PPTA, MPTA, and seven EPTA backends use the parameters $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and $\mathrm{ECORR}$} [39, Eq. 2]. InPTA and the remaining EPTA backends only use $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [43].
@@
Table 1: Priors for the WN parameters, and for the amplitudes/exponents of the power-law RN, DM and SGWB noise spectra. “S” denotes single pulsar analysis; “F” denotes multipulsar analysis. The WN dictionary found in the “S” stage is used in the “F” stage.
The WN parameters for a given pulsar are obtained from all of its (.tim files) TOA data. These are analyzed using JAXopt [44], with priors taken from the Table 1 rows marked “S”; the amplitude/exponent priors are used for all power laws. For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model, while for $\gamma>1$ the $f\to\infty$ divergences are absent [45, 46]. Note that $\mathrm{EQUAD}$} has the units used in Tempo [47], Tempo2 [41], and PINT [48], and is written as $\rm T2EQUAD$. The output is a “WN dictionary” of best-fit $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (potentially) $\mathrm{ECORR}$} values, used later to obtain $\mathbf{N}$} [39, Eq. 2].
For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values. This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters. In SM Sec. I, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
@@
Within the Discovery framework, we use the WN dictionary plus the SGWB priors listed in Tab. 1 with an “F” in the final column. As described above, we consider two priors for the SGWB power-law power-spectrum exponent $\gamma_{gw}$. One is the uniform range given in Tab. 1, the other is fixed to $\gamma_{gw}=13/3$. Our analysis uses the Hamiltonian No-U-Turn Sampler (NUTS) [50, 51], as implemented in NumPyro [52, 53, 54]. NUTS is the default for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers. It generates samples that are much more independent than the parallel tempering MCMC sampler [55] previously used by the individual PTAs, and is also much faster.
We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number under the CURN hypothesis. Each gives the amplitudes and exponents of the power-law models describing the pulsar (red and DM) noise, and the SGWB. We use these to evaluate the frequentist detection statistics and their corresponding false-alarm probabilities ($p$-values) for each posterior sample.
@@
Results of the five PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1). These are shown in Fig. 2. The amplitude posterior is sharply peaked away from zero, which is consistent with (or evidence for) the presence of an SGWB. As expected, the posteriors are very similar for the two hypotheses. This is because, if there are $\lesssim 1500$ pulsars (see [56, Fig. 10]), most of the SGWB information is provided by auto- rather than cross-correlations, and by construction, CURN includes these autocorrelations.
@@
CURN with free $\gamma_{gw}\qquad$ $14.37^{+0.11}_{-0.12}\qquad$ $3.63^{+0.27}_{-0.26}$